OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − 3 = x2 + 4x?

OpenStudy (campbell_st):

rewrite the equation as \[y = x^2 + 4x + 3\] can you factorise this..?

OpenStudy (campbell_st):

to find the x- intercepts let y = 0 then \[x^2 + 4x + 3 = 0\]

OpenStudy (anonymous):

but like isnt it a rule that a cant equal 0?

OpenStudy (campbell_st):

the rule for x intercepts is that y = 0.... you are finding the points on the x-axis.... this is where y is always equal to zero

OpenStudy (anonymous):

no. A cant equal zero... and it does in this equation

OpenStudy (campbell_st):

you let the equation equal zero \[x^2 + 4x + 3 = 0\] so can you factorise

OpenStudy (anonymous):

@marmar3000 , if you're referring to the coefficients in the equation \(\large y=ax^2+bx+c \), then in your equation a is not zero...

OpenStudy (campbell_st):

now you need to find the factors of 3 that add to 4